An object of mass 2 kg is launched at an angle of 30o above the ground with an initial speed of 40 m/s. Neglecting air resistance , calculate:

i. the kinetic energy of the object when it is launched from the  the ground.

 

ii. the maximum height attained by the object .

 

iii. the speed of the object when it is 12 m above the ground.

 


1
Expert's answer
2022-06-13T08:23:14-0400

Given:

"m=2\:\rm kg"

"\theta=30^\circ"

"v_0=40\:\rm m/s"

"g=9.8\:\rm m/s^2"


(i)

"K_0=\frac{mv_0^2}{2}=\frac{2*40^2}{2}=1600\:\rm J"

(ii)

"h_{\max}=\frac{(v_0\sin\theta)^2}{2g}=\frac{(40\sin30^\circ)^2}{2*9.8}=20\:\rm m"

(iii)

"v=\sqrt{v_0^2-2gh}=\sqrt{40^2-2*9.8*12}=37\:\rm m/s"